# LibMOON: A Gradient-based MultiObjective OptimizatioN Library in PyTorch

### v1

@@ Abstract
Multiobjective optimization problems (MOPs) are prevalent in machine learning, with applications in multi-task learning, learning under fairness or robustness constraints, etc. Instead of reducing multiple objective functions into a scalar objective, MOPs aim to optimize for the so-called Pareto optimality or Pareto set learning, which involves optimizing more than one objective function simultaneously, over models with millions of parameters. Existing benchmark libraries for MOPs mainly focus on evolutionary algorithms, most of which are zeroth-order methods that do not effectively utilize higher-order information from objectives and cannot scale to large-scale models with millions of parameters. In light of the above gap, this paper introduces LibMOON, the first multiobjective optimization library that supports state-of-the-art gradient-based methods, provides a fair benchmark, and is open-sourced for the community.
@@ 1 Introduction
$\min_{\theta}{\bm{L}}(\theta)=(L_{1}(\theta),\ldots,L_{m}(\theta)),$ (1)
where $L_{1}(\theta),\ldots,L_{m}(\theta)$ denote $m$ (potentially) conflicting objectives and we denote the size of the model parameter as $N\coloneqq|\theta|$. Note that as informally defined above, Equation 1 is a vector optimization problem that does not necessarily admit a total ordering. Nevertheless, we use the formulation above to emphasize that the underlying problem involves $m$ objectives and we will formally define the goal as follows.
For non-trival multiobjective problems, no single solution can attain the minimum of all objectives simultaneously. To compare two solutions for an MOP, we introduce the concepts of dominance and Pareto optimality [8]. Dominance occurs when solution $\theta^{(a)}$ satisfies $L_{i}(\theta^{(a)})\leq L_{i}(\theta^{(b)})$ for all $1\leq i\leq m$, with at least one strict inequality $L_{j}(\theta^{(a)})<L_{j}(\theta^{(b)})$. A solution is Pareto optimal if no other solution in the feasible region dominates it. The set of all Pareto optimal solutions is the Pareto set (PS), and its image is the Pareto front (PF).
@@
Compared to MOEAs, gradient-based multiobjective optimization (MOO) methods are particularly suitable designed for large-scale machine learning tasks involving thousands to millions of neural network parameters. While these methods can find Pareto stationary solutions—solutions that cannot be locally improved in any objective—they are also effective approximators of Pareto solutions for nonconvex models, including deep neural networks.
In this paper, we roughly divide recent gradient-based MOO methods into two main groups. The first one focuses on finding a finite set of Pareto solutions to approximately represent the entire Pareto front, exemplified by the Upper Bounded Multiple Gradient Descent Algorithm (MGDA-UB) [12] and Exact Pareto Optimization (EPO) [13]. The second, known as Pareto set learning (PSL) [14, 15, 16], develops a single neural model parameterized by $\phi$ to represent the entire Pareto set. Since the parameter space of $\phi$ is large, thus PSL is usually optimized by gradient-based methods.
@@
In this paper, bold letters represent vectors (e.g., ${\bm{\lambda}}$ for preference vectors, ${\bm{L}}(\cdot)$ for vector objective functions), while non-bold letters represent scalars or network parameters (e.g., $\theta$). The number of objectives is denoted by $m$, and $K$ represents the size of finite solution set. The preference vector ${\bm{\lambda}}$ lies in the $(m-1)$-dimensional simplex, satisfying $\sum_{i=1}^{m}\lambda_{i}=1$ and $\lambda_{i}\geq 0$. The decision network parameter $\theta$ has a size of $n$. Refer to Table 12 for the full notation table.
@@ 2.1 Gradient-based multiobjective optimization
Gradient-based MOO has a long research history. For instance, the famous convex optimization book [22]Chap 4 discusses using linear scalarization to convert MOO into a single-objective optimization (SOO) problem. Historically, gradient-based methods were not mainstream for multi-objective optimization, with evolutionary algorithms (MOEAs) being more prevalent due to their ability to maintain a population, aligning well with the Pareto set. However, gradient-based MOO has gained popularity recently with advancements in machine learning. A seminal work is MGDA-UB [12], which introduced MOO concepts into deep learning by framing MTL as MOO. Since then, numerous approaches like EPO [13], Pareto MultiTask learning (PMTL) [17], and MOO with Stein Variational Gradient Descent (SVGD) [18], including methods for learning the entire Pareto set [14, 15, 23, 24, 25], have emerged. This paper aims to integrate these developments into a unified framework.
@@ 2.2 Multiobjective optimization libraries
There a number of multiobjective library before our work. The difference between LibMOON and previous libraries is that LibMOONis designed to support gradient-based MOO methods. Comparison can be found in Table 1.
LibMTL. [26] LibMTL is a Python-based library for multitask learning. The key difference between LibMOONand LibMTL is that LibMTL focuses on finding a single network to benefit all tasks, such as optimizing direction or network architecture. In contrast, LibMOON addresses inherent trade-offs, where improving one objective inevitably worsens others, and explores the distribution of Pareto solutions.
jMetal. [27] jMetal is a comprehensive Java framework designed for multi-objective optimization, leveraging the power of metaheuristics. It offers a highly flexible and extensible platform, making it accessible for users across various domains. With its user-friendly interface and robust architecture, jMetal has become a popular choice for researchers and practitioners in diverse application areas.
Pymoo. [28] Pymoo is a Python-based framework for multiobjective optimization problems with many local optima. It supports mainstream MOEA methods such as NSGA-II [29], NSGA-III [30, 31], MOEA/D [32], and SMS-EMOA [33]. Pymoo allows flexible algorithm customization with user-defined operators and includes tools for exploratory analysis and data visualization. It also offers multi-criteria decision-making strategies to help users select optimal solutions from the Pareto set.
@@
PlatEMO Matlab 2017 × EC
@@ 3 LibMOON: A gradient-based MOO library in PyTorch
To introduce LibMOON , we first introduce its framework in Section 3.1, and briefly introducing its supporting problems and metrics. Then introduce supported solvers in Sections 3.2, 3.3 and 3.4.
@@ 3.1 Framework
Figure 1 demonstrates the framework of the proposed LibMOON library. LibMOON supports three categories of solvers: MOO solvers aiming to find a finite set of Pareto solutions satisfying certain requirements, PSL solvers aiming to learn whole PS with a single model, and MOBO solvers aiming to solve expensive problems. Each solver category is designed highly modulized and new solvers are easy to plugin LibMOON by rewriting only a small portion of code, e.g., the way of gradient manipulation  [^1] An example of adding a new method is given in https://libmoondocs.readthedocs.io/en/latest/develop/add_method.html.. MOO and PSL solvers support all synthetic, MTL, and realworld (RE) problems, while MOBO solvers support synthetic and RE problems.
@@
LibMOON supports a wide range of metrics namely, (1) hypervolume (HV), (2) inverted general distance (IGD), (3) fill distance (FD), minimal distance ($\mathrm{l}_{\min}$), (4) smooth minimal distance ($\mathrm{sl}_{\min}$), (5) Spacing, (6) Span, (7) penalty-based intersection (PBI), (8) inner product (IP), (9) cross angle ($\vartheta$). Full descriptions are provided in Table 13.
@@ 3.2 MOO solvers
To find Pareto solutions, a straightforward way is to use an aggregation function $g_{\lambda}^{\mathrm{agg}}:\mathbb{R}^{m}\to\mathbb{R}$ to transform a multi-objective problem (MOP) into a single-objective optimization problem (SOP). This involves minimizing:
$\min_{\theta}g_{\lambda}^{\mathrm{agg}}({\bm{L}}(\theta)),$ (2)
where ${\bm{L}}(\theta)$ represents the objective vector and $g_{\lambda}^{\mathrm{agg}}$ is an aggregation function determined by preference $\lambda$. If $g_{\lambda}^{\mathrm{agg}}$ is decreasing with respect to ${\bm{L}}(\theta)$ (i.e., $g_{\lambda}^{\mathrm{agg}}({\bm{L}}(\theta^{(a)}))<g_{\lambda}^{\mathrm{agg}}({\bm{L}}(\theta^{(b)}))$ when ${\bm{L}}_{i}(\theta^{(a)})\leq{\bm{L}}_{i}(\theta^{(b)})$ for all $i$ with at least one strict inequality), the optimal solution of Equation 2 corresponds to a Pareto optimal solution of the original MOP (Equation 1) [41].
The linear scalarization (LS) function, $g_{\lambda}^{\mathrm{agg}}({\bm{L}}(\theta))=\sum_{i=1}^{m}\lambda_{i}L_{i}(\theta)$, is the most commonly used aggregation function. Other notable functions include the (Smooth) Tchebycheff (Tche) [24, 41], modified Tchebycheff (mTche) [42], Penalty-Based Intersection (PBI) [32], and COSMOS [36]. For detailed formulations, see Appendix A. Aggregation-based methods can be efficiently optimized in LibMOON using single objective backpropagation. Aggregation-based method update the parameter $\theta$ iteratively as follows:
$\theta\xleftarrow{}\theta-\eta{\bm{d}}=\theta-\eta\frac{\partial g_{\lambda}^{\mathrm{agg}}({\bm{L}}(\theta))}{\partial\theta},$ (3)
where ${\bm{d}}$ is the updating direction.
@@
In addition to optimizing the single-objective aggregation function, multiple gradient manipulation methods adjust gradients and compute update directions to ensure solutions satidfy specific properties. These methods can be summarized as follows: the update direction ${\bm{d}}$ is a dynamically weighted sum of gradient vectors [17], expressed as ${\bm{d}}=\sum_{i=1}^{m}\tilde{\alpha}_{i}\cdot\nabla L_{i}(\theta)$. Equivalently, updating the direction involves optimizing a generalized aggregation function.
$\tilde{g}_{\lambda}({\bm{L}}(\theta))=\sum_{i=1}^{m}\tilde{\alpha}_{i}\cdot L_{i}(\theta).$ (4)
@@
As listed in Table 3, notable methods include EPO [13] for finding ‘exact Pareto solutions’ (intersection points of Pareto front and preference vectors), HVGrad [38] for maximizing hypervolume, MGDA-UB [12] for finding directions beneficial to all objectives, MOO-SVGD [18] for diverse solutions, PMGDA [39] for Pareto solutions satisfying specific user requirement, and PMTL [17] for sector-constrained Pareto solutions.
In LibMOON , multiple gradient manipulation methods are implemented by first calculating dynamic weight factors $\tilde{\alpha}_{i}$ and then performing backpropagation on the generalized function $\tilde{g}_{\lambda}({\bm{L}}(\theta))$ to update the neural network parameters. Those weight factors $\tilde{\alpha}_{i}$ are typically computed by solving a linear programming (LP) problem (e.g., [13][Eq. 24], [39][Eq. 16]), a quadratic programming (QP) problem (e.g., [17][Eq. 14], [12][Eq. 3]), or through other methods like hypervolume gradient estimation [43].
@@
LibMOON not only supports first-order optimization when the gradients of those objective functions are known, but also supports zero-order optimization methods with estimated gradients ^$\hat{\nabla}L_{i}(\theta)$ using methods such evolutionary strategy (ES) [44].
@@ 3.3 Pareto set learning solvers
Method Property Paper Gradient $A$ Gradient $B$
@@
Aggregation-based PSL [15] Minimal aggregation function solutions paper BP BP
@@
In addition to the existing methods for finite Pareto solutions, LibMOON supports Pareto Set Learning (PSL), which trains a model with parameter $\phi$ to generate an infinite number of approximate Pareto solutions. We denote the model output as $\theta$ and represent the model as $\theta_{\phi}(\cdot):\Delta_{m}\to\mathbb{R}^{n}$, where $\Delta_{m}$ is the $m$-dim preference simplex and $\mathbb{R}^{n}$ is the entire decision space of $\theta$.
@@
$\min_{\phi}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}(\Delta_{m})}\tilde{g}_{\bm{\lambda}}({\bm{L}}(\theta_{\phi}({\bm{\lambda}}))),\qquad\theta_{\phi}(\cdot):\Delta_{m}\xrightarrow{}\mathbb{R}^{n}.$ (5)
The function $\tilde{g}_{\bm{\lambda}}(\cdot):\mathbb{R}^{m}\to\mathbb{R}$ is a generalized aggregation function (as introduced in Equation 4), which converts an $m$-dimensional objective vector into a scalar based on specific preferences. Since Equation 5 involves preference expectation, all preference-based finite Pareto solvers (Table 3) can serve as basic Pareto solution solvers, while preference-free solvers are unsuitable for PSL.
@@
For PSL, we explain the function ${\bm{L}}(\theta_{\phi}({\bm{\lambda}}))$ with more details as follows: In synthetic problems, the decision variable is directly the model output $\theta_{\phi}$, and $\theta_{\phi}$ is evaluated by the loss function. In multi-task learning (MTL) problems, $\theta_{\phi}({\bm{\lambda}})$ represents the parameters of a target network, and ${\bm{L}}(\theta_{\phi}({\bm{\lambda}}))$ is evaluated using the dataset $(x_{i},y_{i})$ with this target network. See Figure 2 for a visualization of these architectures.
@@
$\underbrace{\frac{\partial\ell_{\mathrm{psl}}}{\partial\phi}}_{1\times D}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}(\Delta_{m})}\underbrace{\frac{\partial\tilde{g}_{\bm{\lambda}}}{\partial{\bm{L}}}}_{A:(1\times m)}\cdot\underbrace{\frac{\partial{\bm{L}}}{\partial\theta}}_{B:(m\times n)}\underbrace{\frac{\partial\theta}{\partial\phi}}_{C:(n\times D)}.$ (6)
The expectation is empirically estimated using a batch of $K$ preferences. The gradient vector $A=(\tilde{\alpha}_{1},\ldots,\tilde{\alpha}_{m})$ is computed by the finite Pareto solution solver. The gradient matrix $C$ can be calculated either via backpropagation (when gradients are easily obtained) or using a zero-order method. In Equation 6, $C$ is always estimated through backpropagation, since $\theta$ is a continuous function of $\phi$. Gradient calculation by existing PSL methods is summarized in Table 4. After calculating the gradient of $\ell_{\mathrm{psl}}$, the parameter $\phi$ is iteratively updated as $\phi\xleftarrow{}\phi-\eta\frac{\partial\ell_{\mathrm{psl}}}{\partial\phi}$.

### v2

@@ Abstract
Multiobjective optimization problems (MOPs) are prevalent in machine learning, with applications in multi-task learning, learning under fairness or robustness constraints, etc. Instead of reducing multiple objective functions into a scalar objective, MOPs aim to optimize for the so-called Pareto optimality or Pareto set learning, which involves optimizing more than one objective function simultaneously, over models with thousands / millions of parameters. Existing benchmark libraries for MOPs mainly focus on evolutionary algorithms, most of which are zeroth-order / meta-heuristic methods that do not effectively utilize higher-order information from objectives and cannot scale to large-scale models with thousands / millions of parameters. In light of the above gap, this paper introduces LibMOON, the first multiobjective optimization library that supports state-of-the-art gradient-based methods, provides a fair benchmark, and is open-sourced for the community.
@@ 1 Introduction
$\min_{\theta}{\bm{L}}({\bm{\theta}})=(L_{1}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})),$ (1)
where $L_{1}({\bm{\theta}}),\ldots,L_{m}({\bm{\theta}})$ denote $m$ (potentially) conflicting objectives and we denote the size of the model parameter as $N\coloneqq|{\bm{\theta}}|$. Note that as informally defined above, Equation 1 is a vector optimization problem that does not necessarily admit a total ordering.
For non-trivial multiobjective problems, no single solution can attain the minimum of all objectives simultaneously. To compare two solutions for an MOP, we introduce the concepts of dominance and Pareto optimality [8]. Dominance occurs when solution ${\bm{\theta}}^{(a)}$ satisfies $L_{i}({\bm{\theta}}^{(a)})\leq L_{i}({\bm{\theta}}^{(b)})$ for all $1\leq i\leq m$, with at least one strict inequality $L_{j}({\bm{\theta}}^{(a)})<L_{j}({\bm{\theta}}^{(b)})$. A solution is Pareto optimal if no other solution in the feasible region dominates it. The set of all Pareto optimal solutions is the Pareto set (PS), and its image is the Pareto front (PF).
@@
Compared to MOEAs, gradient-based multiobjective optimization (MOO) methods are particularly suitable designed for large-scale machine learning tasks involving thousands to millions of neural network parameters. While these methods can find Pareto stationary solutions—solutions that cannot be locally improved in any objective—they are also effective approximators of Pareto solutions for nonconvex models, including deep neural networks [12].
In this paper, we roughly divide recent gradient-based MOO methods into two main groups. The first one focuses on finding a finite set of Pareto solutions to approximately represent the entire PF. Examples include Upper Bounded Multiple Gradient Descent Algorithm (MGDA-UB) [13] and Exact Pareto Optimization (EPO) [14]. The second, known as Pareto set learning (PSL) [15, 16, 17], develops a single neural model parameterized by ${\bm{\phi}}$ to represent the entire Pareto set. Since the parameter space of ${\bm{\phi}}$ is large, thus PSL is usually optimized by gradient-based methods.
@@
In this paper, bold letters represent vectors (e.g., ${\bm{\lambda}}$ for preference vectors, ${\bm{L}}()$ for vector objective functions), while non-bold letters represent scalars. The number of objectives is denoted by $m$, and $K$ represents the size of finite solution set. The preference vector ${\bm{\lambda}}$ lies in the $m$-dim simplex (${\bm{\Delta}}_{m}$), satisfying $\sum_{i=1}^{m}\lambda_{i}=1$ and $\lambda_{i}\geq 0$. The decision network parameter ${\bm{\theta}}$ has a size of $n$. Refer to Table 12 for the full notation table.
@@ 2.1 Gradient-based multiobjective optimization
Gradient-based MOO has a long research history. For instance, the famous convex optimization book [23][Chap 4] discusses how to use linear scalarization to convert MOO into a single-objective optimization (SOO) problem. However, in the last several decades, gradient-based methods were not a mainstream for MOO, with evolutionary algorithms (MOEAs) being more prevalent due to their ability to maintain a population, which is inherently suitable to approximate the Pareto set and avoid local optimal solutions. However, gradient-based MOO has gained growing popularity very recently with their applications in (deep) machine learning. A seminal work in this field is MGDA-UB [13], which introduced MOO concepts into deep learning by framing an MTL problem as an MOO. Since then, numerous approaches such as EPO [14], Pareto MultiTask learning (PMTL) [18], MOO with Stein Variational Gradient Descent (MOO-SVGD) [19], and methods for learning the entire Pareto set [15, 16, 24, 25, 26], have emerged. This paper categorized these methods by two main classes and implement these methods in a standard manner.
@@ 2.2 Multiobjective optimization libraries
A number of multiobjective libraries exist before our work. The difference between LibMOON and previous libraries is that LibMOON is designed to support gradient-based MOO methods. Comparisons with previous is provided in Table 1.
LibMTL [27]. LibMTL is a Python-based library for multitask learning. The key difference between LibMOON and LibMTL is that LibMTL focuses on finding a single network to benefit all tasks, such as calculating a benign updating direction or optimizing a network architecture. In contrast, LibMOON addresses inherent trade-offs in machine learning problems, where improving one objective inevitably worsens others, and explores the distribution of Pareto solutions.
jMetal [28]. jMetal is a comprehensive Java framework designed for multi-objective optimization, leveraging the power of metaheuristics. It offers a highly flexible and extensible platform, making it accessible for users across various domains. With its user-friendly interface and robust architecture, jMetal has become a popular choice for researchers and practitioners in diverse application areas.
Pymoo [29]. Pymoo is a Python-based framework for multiobjective optimization problems with many local optima. Pymoo supports mainstream MOEA methods such as NSGA-II [30], NSGA-III [31, 32], MOEA/D [33], and SMS-EMOA [34]. Pymoo allows flexible algorithm customization with user-defined operators and includes tools for exploratory analysis and data visualization. It also offers multi-criteria decision-making strategies to help users select optimal solutions from the Pareto set.
@@
PlatEMO Matlab 2017 × EC and meta-heuristics
@@ 3 LibMOON: A gradient-based MOO library in PyTorch
This section introduces LibMOON . We first introduce its framework in Section 3.1, and briefly introducing its supporting problems and metrics. Then we introduce supported solvers in Sections 3.2, 3.3 and 3.4.
@@ 3.1 Framework
Figure 1 demonstrates the framework of the proposed LibMOON library. LibMOON supports three categories of solvers: MOO solvers aiming to find a finite set of Pareto solutions satisfying certain requirements, PSL solvers aiming to learn whole PS with a single model, and MOBO solvers aiming to solve expensive problems. Each solver category is designed highly modulized and new solvers are easy to plugin LibMOON by rewriting only a small portion of code, e.g., the specific way of gradient manipulations  [^1] An example of adding a new method is given by https://libmoondocs.readthedocs.io/en/latest/develop/add_method.html.. MOO and PSL solvers support all synthetic, MTL, and realworld (RE) problems, while MOBO solvers support synthetic and RE problems.
@@
LibMOON supports a wide range of metrics namely, (1) hypervolume (HV), (2) inverted general distance (IGD), (3) fill distance (FD), (4) minimal distance ($\mathrm{l}_{\min}$), (5) smooth minimal distance ($\mathrm{sl}_{\min}$), (5) Spacing, (6) Span, (7) penalty-based intersection (PBI), (8) inner product (IP), (9) cross angle ($\vartheta$). Full descriptions of these indicators are provided in Table 13.
@@ 3.2 MOO solvers
To find Pareto solutions, a straightforward way is to use some aggregation functions $g_{\bm{\lambda}}^{\mathrm{agg}}():\mathbb{R}^{m}\to\mathbb{R}$ to convert a multi-objective problem (MOP) into a single-objective optimization problem (SOP). Aggregation function based methods aim to optimize the following objective function:
$\min_{\bm{\theta}}g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}})),$ (2)
where ${\bm{L}}({\bm{\theta}})$ represents an objective vector and $g_{\bm{\lambda}}^{\mathrm{agg}}()$ is an aggregation function determined by a preference vector ${\bm{\lambda}}$. If $g_{\bm{\lambda}}^{\mathrm{agg}}()$ is decreasing with respect to ${\bm{L}}({\bm{\theta}})$ (i.e., $g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}^{(a)}))<g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}^{(b)}))$ when ${\bm{L}}_{i}({\bm{\theta}}^{(a)})\leq{\bm{L}}_{i}({\bm{\theta}}^{(b)})$ for all $i$ with at least one strict inequality), the optimal solution of Equation 2 corresponds to a Pareto optimal solution of the original MOP (Equation 1[42].
The linear scalarization (LS) function, $g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}))=\sum_{i=1}^{m}\lambda_{i}L_{i}({\bm{\theta}})$, is the most commonly used aggregation function. Other notable functions include the (Smooth) Tchebycheff (Tche) [25, 42], modified Tchebycheff (mTche) [43], Penalty-Based Intersection (PBI) [33], and COSMOS [37]. For detailed formulations, see Appendix A. Aggregation-based methods can be efficiently optimized in LibMOON using single objective backpropagation. Aggregation-based method update the parameter ${\bm{\theta}}$ iteratively as follows:
${\bm{\theta}}\xleftarrow{}{\bm{\theta}}-\eta{\bm{d}}={\bm{\theta}}-\eta\frac{\partial g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}))}{\partial{\bm{\theta}}},$ (3)
where ${\bm{d}}$ is called an updating direction.
@@
In addition to optimizing the single-objective aggregation function, a number of gradient manipulation methods adjust gradients and compute update directions to ensure solutions satisfy specific properties. These methods can be summarized as follows: the update direction ${\bm{d}}$ is a dynamically weighted sum of gradient vectors [18], expressed as ${\bm{d}}=\sum_{i=1}^{m}\tilde{\alpha}_{i}\cdot\nabla L_{i}({\bm{\theta}})$. Equivalently, updating the direction involves optimizing a generalized aggregation function.
$\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))=\sum_{i=1}^{m}\tilde{\alpha}_{i}\cdot L_{i}({\bm{\theta}}).$ (4)
@@
As listed in Table 3, notable methods using gradient manipulations include EPO [14] for finding “exact Pareto solutions” (intersection points of Pareto front and preference vectors), HVGrad [39] for maximizing hypervolume, MGDA-UB [13] for finding directions beneficial to all objectives, MOO-SVGD [19] for diverse solutions, PMGDA [40] for Pareto solutions satisfying specific user requirement, and PMTL [18] for sector-constrained Pareto solutions.
In LibMOON , multiple gradient manipulation methods are implemented by first calculating dynamic weight factors $\tilde{\alpha}_{i}$ and then performing backpropagation on the generalized function $\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))$ to update the parameters of neural network. Those weight factors $\tilde{\alpha}_{i}$ are typically computed by solving a linear programming (LP) problem (e.g., [14][Eq. 24], [40][Eq. 16]), a quadratic programming (QP) problem (e.g., [18][Eq. 14], [13][Eq. 3]), or through other methods like hypervolume gradient estimation [44].
@@
LibMOON not only supports first-order optimization when the gradients of those objective functions are known, but also supports zero-order optimization methods with estimated gradients ^$\hat{\nabla}L_{i}({\bm{\theta}})$ using methods such evolutionary strategy (ES) [45].
@@ 3.3 Pareto set learning solvers
Method Property Paper Term ${\bm{A}}$ Term ${\bm{B}}$
@@
Aggregation-based PSL [16] Minimal aggregation values paper BP BP
@@
In addition to the existing methods for finite Pareto solutions, LibMOON supports Pareto Set Learning (PSL), which trains a model with parameter ${\bm{\phi}}$ to generate an infinite number of approximate Pareto solutions. We denote the model output as ${\bm{\theta}}$ and represent the model as ${\bm{\theta}}_{\bm{\phi}}(\cdot):{\bm{\Delta}}_{m}\to\mathbb{R}^{n}$, where ${\bm{\Delta}}_{m}$ is the $m$-dim preference simplex and $\mathbb{R}^{n}$ is the entire decision space of ${\bm{\theta}}$.
@@
$\min_{\bm{\phi}}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}({\bm{\Delta}}_{m})}\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))),\qquad\theta_{\phi}(\cdot):{\bm{\Delta}}_{m}\xrightarrow{}\mathbb{R}^{n}.$ (5)
The function $\tilde{g}_{\bm{\lambda}}(\cdot):\mathbb{R}^{m}\to\mathbb{R}$ is a generalized aggregation function (as introduced in Equation 4), which converts an $m$-dim objective vector into a scalar based on specific preferences. Since Equation 5 involves preference expectation, all preference-based finite Pareto solvers (Table 3) can serve as basic Pareto solution solvers, while preference-free solvers are unsuitable for PSL.
@@
For PSL, we explain the function ${\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))$ with more details as follows: In synthetic problems, the decision variable is directly the model output ${\bm{\theta}}_{\bm{\phi}}$, and ${\bm{\theta}}_{\bm{\phi}}$ is evaluated by the loss function. In multi-task learning (MTL) problems, ${\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}})$ represents the parameters of a target network, and ${\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))$ is evaluated using the dataset $({\bm{x}}_{i},{\bm{y}}_{i})$ with this target network. See Figure 2 for a visualization of these architectures.
@@
$\underbrace{\frac{\partial\ell_{\mathrm{psl}}}{\partial{\bm{\phi}}}}_{1\times D}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}({\bm{\Delta}}_{m})}\underbrace{\frac{\partial\tilde{g}_{\bm{\lambda}}}{\partial{\bm{L}}}}_{{\bm{A}}:(1\times m)}\cdot\underbrace{\frac{\partial{\bm{L}}}{\partial{\bm{\theta}}}}_{{\bm{B}}:(m\times n)}\underbrace{\frac{\partial{\bm{\theta}}}{\partial{\bm{\phi}}}}_{{\bm{C}}:(n\times D)}.$ (6)
The expectation is empirically estimated using a batch of $K$ preferences. The gradient vector ${\bm{A}}=(\tilde{\alpha}_{1},\ldots,\tilde{\alpha}_{m})$ is computed by the finite Pareto solution solver. The gradient matrix ${\bm{B}}$ can be calculated either via backpropagation (when gradients are easily obtained) or using a zero-order method. In Equation 6, ${\bm{C}}$ is always estimated through backpropagation, since ${\bm{\theta}}$ is a continuous vector function of ${\bm{\phi}}$. Gradient calculation by existing PSL methods is summarized in Table 4. After calculating the gradient of $\ell_{\mathrm{psl}}$, the parameter ${\bm{\phi}}$ is iteratively updated as ${\bm{\phi}}\xleftarrow{}{\bm{\phi}}-\eta\frac{\partial\ell_{\mathrm{psl}}}{\partial{\bm{\phi}}}$.